Activity Expected Cost           Activity Driver Activity Capacity
Setting up equipment $132,000           Number of setups 120
Ordering materials 27,600           Number of orders 1,200
Machining 118,300           Machine hours 9,100
Receiving 34,320           Receiving hours 1,320

Silven produces two models of cell phones with the following expected activity demands:

      Model X       Model Y
Units completed 5,900   11,800
Number of setups 80   40
Number of orders 400   800
Machine hours 5,200   3,900
Receiving hours 440   880

1. Determine the total overhead assigned to each product using the four activity drivers.

Total Overhead Assigned
Model X $
Model Y $

2. Determine the total overhead assigned to each model using the two most expensive activities. The costs of the two relatively inexpensive activities are allocated to the two expensive activities in proportion to their costs. Round interim calculations to two decimal places and round your final answers to the nearest dollar.

Total Overhead Assigned
Model X $
Model Y $

3. Using ABC as the benchmark, calculate the percentage error. Round your answers to two decimal places.

Percentage Error
Model X %
Model Y %
Answer #1

Answer-1- SILVEN COMPANY

Activity Expected Cost Activity Capacity Activity Driver Cost per Activity
Setting up equipment $132,000 120 Number of setups 1,100 per setup
Ordering materials 27,600 1,200 Number of orders 23 per order
Machining 118,300 9,100 Machine hours 13 per machine hour
Receiving 34,320 1,320 Receiving hours 26 per receiving hours
Activity Cost per Activity Model X Model Y
Setting up equipment 1,100 per setup 80 88,000 40 44,000
Ordering materials 23 per order 400 9,200 800 18,400
Machining 13 per machine hour 5,200 67,600 3,900 50,700
Receiving 26 per receiving hour 440 11,440 880 22,880
176,240 135,980

2-Setting up equipment-132,000+[(132,000/ 250,300 (132,000+118,300))*(27,600+34,320)]

=164,655

Machining-118,300+[(118,300/ 250,300)*(27,600+34,320)]

=147,565

New Activity Rates

Setting up equipment- 164,655/ 120= $1,372.13 per setup

Machining- 147,565/ 9,100= $16.22 per machine hour

Model X Model Y
Setting up equipment
1,372.13*80 109,770
1,372.13*40 54,885
Machining
16.22*5,200 84,344
16.22*3,900 63,258
Total overhead assigned 194,114 118,143

3-Model x- (194,114-176,240)/ 176,240 = 10.14%

Model y- (118,143-135,980)/ 135,980= 13.12%